# Learnable Tokenizer for LLM-based Generative Recommendation

### v1

@@ 4.3.1. Ablation Study (RQ2)
To investigate the effect of each regularization, we compare the following five variants of LETTER on TIGER: (0) We only employ semantic regularization for item tokenization, which is equivalent to TIGER. (1) We incorporate semantic and collaborative regularization for item tokenization, denoted as “TIGER w/ c. r.”. (2) We involve semantic and diversity regularization for item tokenization, referred to as “TIGER w/ d. r.”. (3) We utilize semantic, collaborative, and diversity regularization for item tokenization and train TIGER with original generation loss, denoted as “(1) w/ d. r.”. (4) We employ all regularization for item tokenization and apply ranking-guided generation loss, i.e., LETTER-TIGER.
@@
Results on Instruments and Beauty are depicted in Table 2. We omit the results on Yelp with similar observations to save space. From the table, we have several key observations: 1) Incorporating either collaborative or diversity regularization can improve the performance of TIGER, which validates the effectiveness of injecting collaborative signals into code assignment and improving the diversity of code embeddings, respectively. 2) Optimizing item tokenization with all three regularizations will further improve the performance (superior performance of (3) than (0-2)), indicating the effectiveness of considering both semantics and collaborative information in the code assignment with improved diversity. 3) LETTER-TIGER achieves the best performance, which leverages ranking-guided generation loss. This shows the effectiveness of suppressing hard-negative samples by alterning temperature to strengthen the ranking ability.
@@ 4.3.2. Code Assignment Distribution (RQ2)
From Figure 6, we observe that: 1) incorporating diversity regularization effectively promotes a more uniform distribution of the first code, thus alleviating code assignment bias and potentially mitigating the item generation bias of generative recommender models. 2) The incorporation of diversity regularization significantly raises the utilization rate of codes in the first-level codebook, thereby improving diversity in code assignment. 3) Although adding collaborative regularization decreases the code utilization of TIGER (from 148 to 76), integrating diversity regularization inversely compensates for the drop in code utilization (more than a doubled code utilization rate). As such, LETTER can capture collaborative signals and maintain a high code utilization, simultaneously meeting the multiple criteria of an ideal identifier.
@@ 4.3.3. Code Embedding Distribution (RQ2)
| TIGER | 0.0050 | 0.0150 | 0.0024 | 0.0049 | |
| Instruments | LETTER | 0.0080 | 0.0159 | 0.0038 | 0.0058 |
@@
Similarity experiment. We aim to verify whether the items with similar collaborative signals exhibit similar identifiers. We first pair every item with its most similar item based on the similarity from pre-trained CF embeddings. Next, we assess the similarity of the code sequence between the two items within the item pairs. Specifically, we measure the overlap degree between the code sequences of the two items and report the averaged results of all items in Table 4. We can find that LETTER achieves more similar code sequences for the items with similar collaborative signals than TIGER, thereby alleviating the misalignment issue between semantics and the collaborative similarity.
| Beauty | Instruments | |
| TIGER | 0.1135 | 0.0849 |
| LETTER | 0.3312 | 0.2760 |
@@ 4.3.5. Hyper-Parameter Analysis (RQ3)
Identifier length . We vary from 2 to 8 and present the results in Figure 8. It is observed that 1) the performance improves as we increase from 2 to 4. This is because overly short identifiers may lose some fine-grained information, resulting in weaker expressiveness of idetentifiers. 2) Continuously increasing the identifier length from 4 to 8 will inversely degrade the performance. This is reasonable because autoregressive generation suffers from error accumulation (Ren et al., 2024b). Since accurate item generation requires the correctness of all generated codes in the sequence, generating longer identifiers accurately is more challenging than shorter identifiers.
Codebook size . We compare LETTER with 64, 128, 256, and 512. Results in Figure 8 show that 1) a small codebook size (N=64) yields inferior performance while gradually increasing tends to perform better. The worse performance of small-sized codebooks may be caused by the limited diversity of code selection, thus failing to distinguish items effectively. However, 2) blindly expanding might inversely hurt the performance (inferior performance when ). We speculate that larger codebooks are susceptible to noises in the item’s semantic information, potentially leading to the over-fitting of some meaningless semantics and thus hindering accurate recommendation.
@@
Strength of diversity regularization . We examine the influence of different diversity regularization strengths by adjusting from 0.00001 to 0.01. The results are in Figure 8 and we observe that applying diversity regularization with a slight strength is sufficient to enhance the diversity of code assignments (significant improvements from to ). In contrast, an excessive amount of diversity signal may interfere with the tokenizer’s integration of semantic and collaborative signals.
Temperature . From the results of LETTER-TIGER with multiple values of ranging from 0.6 to 1.2 as shown in Figure 8, we can find that decreasing from 1.2 to 0.7, the performance tends to incline. This is reasonable as decreasing temperature casts more emphasis on the penalty for hard negatives, where the ranking ability is strengthened, thus improving the recommendation performance. Nevertheless, we should carefully choose since a too-small might suppress the possibility of hard negative samples being considered as positive samples for other users.
@@ 5. Related Work
LLMs for Generative Recommendation. LLMs have shown promising prospects for generative recommendation (Lin et al., 2023; Gong et al., 2023), where item tokenization is a crucial problem. Existing work on item tokenization primarily investigate three types of identifiers for item indexing: ID identifiers (Hua et al., 2023; Geng et al., 2022; Wang et al., 2024; Chu et al., 2023), textual identifiers (Bao et al., 2023a; Zhang et al., 2021, 2023a; Li et al., 2023; Liao et al., 2024; Dai et al., 2023), codebook-based identifiers (Wang et al., 2024; Rajput et al., 2023; Zheng et al., 2024). In early stages, randomly assigned IDs struggled to effectively encode the semantic information and collaborative patterns of items. So ID identifiers such as SemID (Hua et al., 2023) and CID (Hua et al., 2023) use item content semantic information (i.e., attributes) and collaborative signals from item co-occurrences, respectively, to build tree-like structures for generating identifiers. However, such a fixed and unlearnable structure is hard to efficiently and effectively represent item similarity and adapt new items. Alternatively, textual identifiers leverage detailed descriptive information of items as identifiers. Nevertheless, these methods suffer from length bias and struggle to hierarchically encode semantic information or integrate collaborative signals. Codebook-based identifiers leverage codebooks and attempt to integrate semantic and collaborative signals during training. However, they overlook the misalignment introduced by incorporating CF information into fixed code sequences, as well as the code assignment bias. In this work, we propose a superior identifier that seamlessly integrates hierarchical semantic with collaborative signals and code assignment diversity.
@@ 6. Conclusion and Future Work
This work highlights the importance of item tokenization on generative recommendation, leaving some promising direction for future exploration. 1) Tokenization with user behaviors: this work only considers user liking behaviors in the user historical interactions. In the future, we can explore tokenization with rich user behaviors to enable generative recommender models to infer user preference from multiple user actions. 2) Cross-domain item tokenization: LETTER has the potential to tokenize cross-domain items for open-ended recommendation, enabling generative recommender models to leverage multi-domain user behaviors and items for user preference reasoning and next-item recommendation. 3) Integration of user instructions and tokenized interaction history for generative recommendation: it is promising to combine user instructions in natural language with user interaction history tokenized by LETTER for personalized recommendations, achieving collaborative reasoning with complex natural language instructions and item tokens in the space of generative recommender models.
@@ 7. Appendix
Our proof is primarily divided into two parts. The first part establishes the effectiveness of our loss function in mining hard negative samples. The second part demonstrates the correlation between our loss function and ranking.
7.0.1. Hard Negative Mining
This subsection primarily proofs two objects: Firstly, the ranking-enhanced loss can mine hard negative samples. Secondly, as the value of decreases, there is an increased focus on hard negative samples. We first prove hard negative mining ability of :
Next, we shall demonstrate that when t-1 tokens are determined, our loss is minimized by performing hard negative mining on the t-th token.
Next, we analyze the gradient associated with hard negative mining. It is worth noting that this gradient is equivalent to minimizing the hard negative mining loss as follows:
where denotes stop-gradient operation. The key of hard negative mining lies in . We can observe that:
This means the harder the negative sample is, the larger is. Such observation demonstrate the ability of hard negative sample mining of our ranking-guided loss .
Next, we elucidate that with the diminishing value of , there is a heightened emphasis on hard negative samples. For a given negative sample , we have:
Thus, if
@@ 7.0.1. Hard Negative Mining
This indicates that more difficult negative samples, in other words, negative samples with a higher p(v), have their weights rapidly increasing as decreases. On the other hand, simpler negative samples exhibit the opposite behavior. This process elucidates how our ranking-enhanced loss can modify the difficulty of negative sampling by adjusting the value of .
7.0.2. Relation with Ranking
Subsequently, we reveal the correlation between our loss and ranking metric. Firstly, let us observe the gradient in Eq.10, we have:
@@ 7.0.2. Relation with Ranking
which demonstrates that is a upper bound for every . Similar to (Hu and Hong, 2013), we have minimize is the close form of the following optimization problem:
@@
(Brought from Theorem 2 in (Shi et al., 2023)) optimize problem 18 is associated with optimizing one-way partial AUC (OPAUC). Furthermore, is a surrogate version of OPAUC() object. The formation of OPAUC(), can be presented as following (Zhu et al., 2022):
@@
Subsequently, we shall demonstrate the relation between and top-K ranking metric:

### v2

@@ 4.3.1. Ablation Study (RQ2)
To thoroughly investigate the effect of each regularization, we compare the following five variants of LETTER on TIGER: (0) we solely employ semantic regularization for item tokenization, which is equivalent to TIGER. (1) We incorporate both semantic and collaborative regularization for item tokenization, denoted as “TIGER w/ c. r.”. (2) We involve both semantic and diversity regularization for item tokenization, referred to as “TIGER w/ d. r.”. (3) We utilize semantic, collaborative, and diversity regularization for item tokenization and train TIGER with original generation loss, denoted as “(1) w/ d. r.”. (4) We employ all regularizations for item tokenization and apply ranking-guided generation loss, i.e., LETTER-TIGER.
@@
Results on Instruments and Beauty are depicted in Table 2. We omit the results on Yelp with similar observations to save space. From the table, we have several key observations: 1) incorporating either collaborative or diversity regularization can improve the performance of TIGER, which validates the effectiveness of injecting collaborative signals into code assignment and improving the diversity of code embeddings, respectively. 2) Optimizing item tokenization with all three regularizations will further improve the performance (superior performance of (3) than (0-2)), indicating the effectiveness of considering both semantics and collaborative information in the code assignment with improved diversity. 3) LETTER-TIGER achieves the best performance, which leverages ranking-guided generation loss. This shows the effectiveness of suppressing hard-negative samples by altering temperature to strengthen the ranking ability.
@@ 4.3.2. Code Assignment Distribution (RQ2)
From Figure 6, we observe that: 1) incorporating diversity regularization effectively promotes a more uniform distribution of the first code, thus alleviating code assignment bias and potentially mitigating the item generation bias of generative recommender models. 2) The incorporation of diversity regularization significantly raises the utilization rate of codes in the first-level codebook, thereby improving diversity in code assignment. 3) Although adding collaborative regularization decreases the code utilization of TIGER (from 148 to 76), integrating diversity regularization inversely compensates for the drop in code utilization. As such, LETTER can capture collaborative signals and maintain a high code utilization, simultaneously meeting the multiple criteria of an ideal identifier.
@@ 4.3.3. Code Embedding Distribution (RQ2)
| Instruments | Beauty | |
| TIGER | 0.0849 | 0.1135 |
| LETTER | 0.2760 | 0.3312 |
@@
Similarity experiment. Our target is to verify whether the items with similar collaborative signals exhibit similar identifiers. We first pair every item with its most similar item based on the similarity from pre-trained CF embeddings. Next, we assess the similarity of the code sequence between the two items within the item pairs. Specifically, we measure the overlap degree between the code sequences of the two items and report the averaged results of all items in Table 4. We can find that LETTER achieves more similar code sequences for the items with similar collaborative signals than TIGER, thereby alleviating the misalignment issue between semantics and the collaborative similarity.
@@ 4.3.5. Hyper-Parameter Analysis (RQ3)
Identifier length . We vary from 2 to 8 and present the results in Figure 8. It is observed that 1) the performance improves as we increase from 2 to 4. Because overly short identifiers may lose some fine-grained information, resulting in weaker expressiveness of idetentifiers. 2) Continuously increasing the identifier length from 4 to 8 will inversely degrade the performance. This is reasonable because autoregressive generation suffers from error accumulation (Ren et al., 2024b). Since accurate item generation requires the correctness of all generated codes in the sequence, generating longer identifiers accurately is more challenging than shorter identifiers.
Codebook size . We evaluate LETTER with different codebook sizes 64, 128, 256, and 512. Results in Figure 8 show that 1) Gradually increasing tends to perform better. The worse performance of small-sized codebooks may be caused by the limited diversity of code selection, thus failing to distinguish items effectively. However, 2) blindly expanding might inversely hurt the performance, where larger codebook size might be more susceptible to noise in the item’s semantic information, potentially leading to the over-fitting of some meaningless semantics.
@@
Strength of diversity regularization . We examine the influence of different diversity regularization strengths by adjusting from 0.00001 to 0.01. The results are in Figure 8 and we observe that applying diversity regularization with a slight strength is sufficient to enhance the diversity of code assignment (significant improvements from to ). In contrast, an excessive amount of diversity signal may interfere with the tokenizer’s integration of semantic and collaborative signals.
Cluster . Based on the results with various values of ranging from 5 to 20, as illustrated in Figure 8, we can observe that decreasing or increasing from 10, the performance tends to decline. This is understandable because overly large clusters contain too many code embeddings, causing embeddings within the same cluster to be insufficiently close, whereas overly small clusters have relatively few code embeddings, leading to embeddings within the same cluster being overly close.
Temperature . From the results of LETTER-TIGER with multiple values of ranging from 0.6 to 1.2 as shown in Figure 8, we can find that decreasing from 1.2 to 0.7, the performance tends to decline. This is reasonable as decreasing temperature casts more emphasis on the penalty for hard negatives, where the ranking ability is strengthened, thus improving the recommendation performance. Nevertheless, we should carefully choose since a too-small might suppress the possibility of hard negative samples being considered as positive samples for other users.
@@ 5. Related Work
LLMs for Generative Recommendation. LLMs have shown promising prospects for generative recommendation (Li et al., 2024c; Lin et al., 2024c; Gong et al., 2023; Lin et al., 2024d), where item tokenization is a crucial problem. Existing work on item tokenization primarily investigate three types of identifiers for item indexing: ID identifiers (Hua et al., 2023; Geng et al., 2022; Wang et al., 2024; Chu et al., 2023), textual identifiers (Bao et al., 2023a; Zhang et al., 2021, 2023a; Liao et al., 2024; Dai et al., 2023; Li et al., 2023), codebook-based identifiers (Wang et al., 2024; Rajput et al., 2023; Zheng et al., 2024). In early stages, randomly assigned IDs struggled to effectively encode semantic information and collaborative patterns of items. So ID identifiers such as SemID (Hua et al., 2023) and CID (Hua et al., 2023) use item semantic information and collaborative signals, respectively, to build tree-like structures for generating identifiers. However, such a fixed and unlearnable structure is hard to efficiently and effectively represent item similarity and adapt new items. Alternatively, textual identifiers leverage detailed descriptive information of items as identifiers. Nevertheless, these methods struggle to hierarchically encode semantic information or integrate collaborative signals. Codebook-based identifiers leverage codebooks and attempt to integrate semantic information and collaborative signals during training process. However, they overlook the misalignment introduced by incorporating collaborative signals into fixed code sequences, as well as the code assignment bias. Specifically, LC-Rec uses the Sinkhorn-Knopp Algorithm to fairly consider intra-layer codes through iterative normalization, missing the essence of code assignment: the distribution of code embeddings. In this work, we propose a superior identifier that integrates hierarchical semantic with collaborative signals and code assignment diversity.
@@ 6. Conclusion and Future Work
This work highlights the importance of item tokenization on generative recommendation, leaving some promising direction for future exploration. 1) Tokenization with rich user behaviors to enable generative recommender models to infer user preference from multiple user actions. 2) LETTER has the potential to tokenize cross-domain items for open-ended recommendation, enabling generative recommender models to leverage multi-domain user behaviors and items for user preference reasoning and next-item recommendation. 3) It is promising to combine user instructions in natural language with user interaction history tokenized by LETTER for personalized recommendations, achieving collaborative reasoning with complex natural language instructions and item tokens in the space of generative recommender models.
@@ 7. Appendix
Our proof is primarily divided into hard negative mining and relation with ranking two parts.
Hard Negative Mining. This subsection primarily proofs two objects: Firstly, the ranking-guided loss can mine hard negative samples. Secondly, as the value of decreases, there is an increased focus on hard negative samples. We first prove hard negative mining ability of :
We shall demonstrate that when t-1 tokens are determined, our loss is associated with hard negative mining on the t-th token.
Next, we analyze the gradient that associated with hard negative mining. Noting that this gradient is equivalent to minimizing the hard negative mining loss where , denotes stop-gradient operation. We can observe that: This means the harder the negative sample is, the larger is, demonstrating the ability of hard negative sample mining of our ranking-guided loss .
Next, we elucidate that with the diminishing value of
, there is a heightened emphasis on hard negative samples. For a given negative sample ,
if
@@
showing adjusting the value of in the ranking-guided loss mechanism increases the weights of more difficult negative samples with higher and decreases the weights of simpler negative samples.
Relation with Ranking.
Subsequently, we reveal the correlation between our loss and ranking metric.
Firstly, let us observe the gradient in Eq.9, we have:
@@
which demonstrates that is a upper bound for every . Similar to (Hu and Hong, 2013), minimizing is the close form of the following optimization problem:
@@
(Brought from Theorem 2 in (Shi et al., 2023)) optimize problem 12 is associated with optimizing one-way partial AUC (OPAUC). Furthermore, is a surrogate version of OPAUC() object. The formation of OPAUC(), can be presented as following (Zhu et al., 2022):
@@
Subsequently, we shall demonstrate the relation between and top-K ranking metric, finally showing the relation between ranking-guided loss and ranking metrics: